OpenStudy (anonymous):

solve for x -1 + √(x+5) = 9

OpenStudy (anonymous):

Here would be the steps: 1. Add 1 to both sides 2. Square both sides. 3. Subtract 5 from both sides.

myininaya (myininaya):

\[\sqrt{x+5}=9+1-x\] Combine like terms then square both sides

OpenStudy (mertsj):